### v3

# Mask of truth: model sensitivity to unexpected regions of medical images

Abstract
The development of larger models for medical image analysis has led to increased performance. However, it also affected our ability to explain and validate model decisions. Models can use non-relevant parts of images, also called spurious correlations or shortcuts, to obtain high performance on benchmark datasets but fail in real-world scenarios. In this work, we challenge the capacity of convolutional neural networks (CNN) to classify chest X-rays and eye fundus images while masking out clinically relevant parts of the image. We show that all models trained on the PadChest dataset, irrespective of the masking strategy, are able to obtain an Area Under the Curve (AUC) above random. Moreover, the models trained on full images obtain good performance on images without the region of interest (ROI), even superior to the one obtained on images only containing the ROI. We also reveal a possible spurious correlation in the Chákṣu dataset while the performances are more aligned with the expectation of an unbiased model. We go beyond the performance analysis with the usage of the explainability method SHAP and the analysis of embeddings. We asked a radiology resident to interpret chest X-rays under different masking to complement our findings with clinical knowledge. Our code is available at https://github.com/TheoSourget/MMC_Masking and https://github.com/TheoSourget/MMC_Masking_EyeFundus
This version of the article has been accepted for publication, after peer review but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available online at: http://dx.doi.org/10.1007/s10278-025-01531-5.
Keywords:
Shortcut learning Model robustness Chest X-ray classification Glaucoma classification
1 Introduction
With the increasing demand for imaging examinations and a shortage of clinicians to perform their analyses, the waiting time for patients has deteriorated. In this context, artificial intelligence (AI) has emerged as a possible solution and has become standard in certain applications as more devices received FDA approval [1]. This was possible thanks to the increased performances of AI models, sometimes even reported as outperforming domain experts on some aspects [2, 3]. However, this performance has been achieved using more complex models with less understanding of the reasons for the classification. This leads to biased models relying on non-relevant parts of the images or shortcuts correlated with the condition in the training data. It affects the evaluation of proposed methods resulting in an overestimation of their true capacities and poor performances when tested on external data [4].
While most research is focused on new methodologies improving performances, techniques to detect [5, 6] and mitigate [7, 8] these biases have been proposed. However, they often investigate a single type of bias and rely on prior knowledge or annotation of this bias. Moreover, biases are often datasets specific, there is therefore a need to perform analyses on more recent datasets to discover their potential biases. Finally, studies are generally limited to the effect on performance metrics, not assessing the representation of the data produced by models or including clinicians to evaluate their method. This last aspect seems essential to ensure the clinical relevancy of results when using generative models to create new images or masking out part of the image.
Our contributions, extending previous works on bias detection and aiming for a better understanding of the effect of spurious correlations, are as follows:
  1. 1.
    We evaluate the capacity of a CNN model to classify chest X-ray and eye fundus images with different masking strategies, highlighting the usage of non-clinically relevant parts of the images.
  2. 2.
    We perform a study with a radiology resident to evaluate the possibility of correctly classifying chest X-rays under different masking strategies.
  3. 3.
    We use embedding analysis and an explainability method to better understand which elements are used but also discuss the limitations of such techniques in this setting.
  4. 4.
    We discuss the limitations of our study and present possible extensions.
2 Related Work
Most studies on artificial intelligence and deep learning focus on presenting new methods to improve performances on benchmark datasets. In this context, CNNs have been widely applied to the medical field  [9, 10, 11]. To reduce the need for large training data of CNN models, capsule networks [12] that better model spatial relationships in images have also been applied to medical images [13, 14]. After the breakthrough of transformers on natural language processing tasks and their adaptation to computer vision, several works present approaches using transformer architectures with medical images models [15, 16, 17] sometimes outperforming CNN models. To improve the generalization capacity of models, some studies present foundation models trained with large-scale datasets that can be applied to various downstream tasks [18, 19]. Finally, with the recent progress on models combining multiple types of data (e.g. image and text), some studies also assess vision-language models [20, 21]. In our work, we do not focus on improving the classification metrics but rather on understanding model behaviours and the limitations of such evaluations. We therefore choose a widely used CNN classifier to conduct our analysis.
Recent studies show that datasets can suffer from problems like label errors, biases or shortcuts in datasets which may impact the reliability of the evaluation. In this work, we focus on shortcuts, also called spurious correlations or confounders, leading models to be biased and use irrelevant information to make decisions [22, 23, 24].
Some studies evaluate the impact of a specific shortcut like Jiménez-Sánchez et al. [5] and Oakden-Rayner et al. [6] who show the effect of chest drains on the classification of pneumothorax or Wargnier et al. [25] who show the influence of the brain’s shape in MRI classification. Boland et al. [26] also investigate which layers are impacted by spurious correlation looking at the difference between a model trained with a base non-biased dataset and a biased version of it.
Using generative models, other methods create new or synthetic data with or without the confounders for a better evaluation or to quantify the effect of specific confounders [27, 28]. Weng et al. [27] use a diffusion model guided with an automatically generated mask of the area which should be modified to edit counterfactual images, detect spurious correlation and evaluate its impact. Pérez-García et al. [28] also use a diffusion model for data editing but indicate both which area should be modified and which area should be kept unchanged as well as a text prompt.
Some, more similar to our method, study the performance of models when removing the area of interest. Bissoto et al. [29] perform the analysis of melanoma classification removing the lesion in the image. Hemelings et al. [30] also evaluate the performances of models on eye fundus photographies under different masking strategies showing good results despite removing the optic disc. Both studies only evaluate models on the same type of images (e.g. train and evaluate without the ROI) and did not assess whether clinicians were able to diagnose the images correctly. Haynes et al. [31] assess the capacity of Covid-19 classifiers to correctly classify images without the lungs. They find that models, especially when not pre-trained on related tasks, are still able to distinguish healthy and non-healthy images. Aslani et al. [32] present a pipeline improving COVID-19 detection in chest X-rays. The pipeline includes the segmentation of the lungs and crops the image to centralize the lungs, removing some possible biases on the image’s borders such as texts. Beyond classification models, shortcut learning is also a phenomenon impacting segmentation models as highlighted by Lin et al. [33].
Prior works showed the variety of possible shortcuts, Juodelyte et al. [34] introduce a taxonomy for confounders and evaluate the robustness of CNN pre-trained on ImageNet [35] and RadImageNet [36] for chest X-ray and CT datasets. They show that despite resulting in similar performances, models pre-trained on ImageNet are more impacted by shortcuts than models pre-trained on RadImageNet.
Being able to explain the outcome of a deep-learning model, often seen as a black box, is crucial, especially in a medical context. To this aim, explainability methods are applied to show the clinical relevancy of the image area used by models to classify samples [9, 37]. Following the combination of image and text to improve performances, it is also used to analyse the data and model behaviours. Kim et al. [38] presents a framework with a foundation model to extract concepts in skin lesion images. These concepts can then be used to detect potential biases or be used as the input of a classification network to produce more explainable decisions following the idea of concept bottleneck models [39].
Our study adds to this line of research by analysing more recent datasets in two applications and including a domain expert to strengthen our findings with clinical knowledge. Moreover, our method doesn’t rely on prior annotation of potential shortcuts but on the knowledge of clinically relevant parts of the image.
3 Materials and Methods
3.1 Datasets
3.1.1 Chest X-rays:
Our first task is chest X-ray classification using the PadChest dataset [40] [^1] Accessed from https://bimcv.cipf.es/bimcv-projects/padchest/. We chose this dataset as it is less studied than more popular datasets such as CheXpert [41] or ChestX-ray14 [42]. The dataset contains more than 160,000 images. We removed images of lateral views (marked as "L" in the "Projection" field of the metadata) and images for which the label was null or included "suboptimal study", "exclude" or "unchanged".
The PadChest dataset doesn’t provide segmentation masks for the lungs, we therefore decided to use the CheXmask dataset [43] [^2] Downloaded from https://physionet.org/content/chexmask-cxr-segmentation-data/0.4/. The dataset contains the segmentation masks of five public datasets including PadChest created with a HybridGNet model and an automated quality assessment of the masks using the Reverse Classification Accuracy (RCA) [44]. We only used masks for which the "Dice RCA (Mean)" is above 0.7 as advised by the dataset providers and we removed images for which no masks were available. 93,203 images were kept after this filtering, examples of the data from PadChest and CheXmask can be seen in Fig. 1. We also evaluate out-of-distribution (OOD) performance on 22,004 images from the ChestX-ray14 dataset’s test set [42] [^3] Downloaded from https://nihcc.app.box.com/v/ChestXray-NIHCC.
PadChest is a multi-class and multi-label dataset containing annotations for 174 labels. However, we decided to restrict our study to five classes also present in the ChestX-ray14 dataset: cardiomegaly, pneumonia, atelectasis, pneumothorax and effusion.
3.1.2 Retinal fundus image:
Our second task is the binary classification of glaucoma in colour eye fundus images. We use the Chákṣu dataset [45] [^4] Downloaded from https://figshare.com/articles/dataset/Ch_k_u_A_glaucoma_specific_fundus_image_database/20123135 which contains 1345 images acquired with three different fundus cameras as in Fig. 2. Similar to the choice of PadChest, we chose Chákṣu as it was recently released and therefore less studied than other eye-fundus datasets. The dataset was labelled by five experts and contains, in addition to the glaucoma decision, the segmentation map of the optic disc and optic cup which are the structures used by clinicians to perform the diagnosis. Multiple fusion techniques are available to combine the annotations: using the mean, the median, the majority and the Simultaneous Truth and Performance Level Estimation (STAPLE) algorithm [46]. Dataset providers showed that fusing using the STAPLE algorithm gives the best alignment with the experts, we therefore used the masks from this fusion in our study. We evaluate OOD performance with data from the AIROGS challenge [10] [^5] Downloaded from https://doi.org/10.5281/zenodo.5793241 and, as the test set of the challenge is not accessible, we decided to use the first 18,000 images of the training set.
3.1.3 Pre-processing:
The pre-processing is similar for all datasets. We resized the images to $512\times 512512 × 512$ and normalized them. If needed, we added black padding to an image to preserve the aspect ratio of the original content after the resize.
(a) Chest X-ray from PadChest
(b) Associated mask from CheXmask
Figure 1: Example of chest X-ray from PadChest dataset and associated mask from CheXmask
(a) With Bosch camera
(b) With Forus camera
(c) With Remidio camera
Figure 2: Examples of eye fundus from each camera in the Chákṣu dataset
3.2 Experimental setting, masking of ROI and evaluation
For the PadChest dataset, we split the images into a train and test set with 80/20 proportions. The testing set was already created in the Chákṣu dataset. For both PadChest and Chákṣu, we train a Densenet-121 model [47] pre-trained on imagenet-1k with all weights frozen except the last dense block and the classification head using a 5-fold cross-validation protocol. The final set of hyperparameters for both datasets can be seen in Table 1, all models of the same dataset use the same set of hyperparameters regardless of the masking strategy. Five versions of the dataset are used to train five models:
  1. 1.
    Full images
  2. 2.
    Keeping only the outside of mask boundaries referred to as "No lungs/disc"
  3. 3.
    Keeping only the outside of a bounding box of the mask referred to as "No lungs/disc BB"
  4. 4.
    Keeping only the inside of the precise mask boundaries referred to as "Only lungs/disc"
  5. 5.
    Keeping only the inside a bounding box of the mask referred to as "Only lungs/disc BB"
The usage of a bounding box is important to remove the shape information but also to include or exclude relevant parts like the heart for the cardiomegaly condition. See Fig. 3 for an example of all strategies on a chest X-ray and an eye fundus image. The models are evaluated using the Area Under the Curve (AUC) as this metric quantifies the ability of models to distinguish positive instances regardless of a chosen decision threshold.
Dataset PadChest Chákṣu
Learning rate 1e-5 1e-4
Batch size 32 32
Loss Cross-entropy Weighted cross-entropy
Maximum epochs 250 250
Early stopping 10 epochs, delta 1e-3 10 epochs, delta 1e-3
Data augmentation Random rotation (+/- 45°)
Random horizontal flip (p=0.5)
Brightness modification (factor 0.7,1.1)
Random rotation (+/- 45°)
Random horizontal flip (p=0.5)
Brightness modification (factor 0.7,1.1)
Table 1: Final training hyperparameters for each dataset. The hyperparameters are the same for all masking strategies of a dataset
(a) Chest X-ray from PadChest
(b) Eye fundus from Chákṣu dataset
Figure 3: Example of all masking strategies used in our study. Each strategy is used to train a separate model
As a baseline, we train logistic regression classifiers using tabular data for the Padchest dataset using the same folds as for the image classification models but removing the samples for which at least one of the features was absent. For the chest x-rays, we train the models with the patient’s birth year, the patient’s sex and the projection view. No patient metadata was available for the eye fundus dataset to conduct such experiments. Baseline models on tabular data are recommended to evaluate the additive value of images compared to metadata [48].
We evaluated the statistical significance of the differences between the AUCs of the different models with the Delong test [49, 50]. The Delong test is the most appropriate test to compare correlated AUC and was for example used in [51, 52] to compare classifiers on medical images. As we have five models per version of the dataset, we concluded that a model was significantly better if the p-values were below 0.05 for at least three folds.
3.3 Comparison of embeddings
In addition to the results of the classification task, we analyse the generated embeddings to evaluate their similarity. We use the embedding vectors of dimension 1024 of the flattened global average pooling layer before the classification head of the model trained with full images applied to images with different masking. We use the t-SNE algorithm [53] to have a global view of the proximity of the different embeddings. The t-SNE algorithm transforms high-dimensional data into lower-dimension (often two or three dimensions to visualize the data) while preserving the local structure of the data in high-dimension. We also compute the cosine similarity between the embeddings of full images and their different masked variation for a quantitative measure.
3.4 Explainability method
We use SHAP (SHapley Additive exPlanations) [54] which divides the image into superpixels and computes the fair contribution of each superpixel using the change in the classifier outputs after their alteration using occlusion, blurring or inpainting method. We used an occlusion mask and set the number of evaluations to 1000.While other works pointed some limitations to its usage [55, 56], we decided to use SHAP following the results of Sun et al. [37]. They show its better performance on medical images containing shortcuts compared to gradient-based explanation methods such as Grad-Cam [57].
3.5 Study with a domain expert
We compare the ability of the AI models and a domain expert to distinguish healthy and non-healthy images to assess whether it was possible to make the correct diagnosis for clinically relevant reasons in the images. It is for example useful for the effusion class which appears in the pleural space and could therefore be visible in images without lungs depending on the masks. We asked a radiology resident with five years of experience and who has been involved previously in multiple AI projects to annotate images from each condition and the different masking strategies. The study was performed in two steps. During the first step, the radiology resident had to annotate ten images randomly selected from our training set and could ask any question. This phase was used to present the study to the radiology resident, to ask his opinion on the annotation tool and to improve it.
During the second step, we selected three images per condition and masking (75 images in total) in the testing set. We selected the images based on the output of the model: one image from the highest probabilities output, one near the median and one among the low probabilities. We informed the radiology resident that an image could contain one or multiple of the conditions but also conditions we do not track or not contain conditions at all. The radiology resident did not have access to other information such as the projection (Posterior-Anterior (PA) and Anterior-Posterior (AP)) which is important for some conditions such as cardiomegaly or pleural effusion as it impacts the visual aspect.
4 Results
4.1 Different impact between chest X-rays and eye fundus
We observed different results between chest X-rays and eye fundus. For example, Fig. 4 shows the AUC for the effusion class of the PadChest dataset and glaucoma of the Chákṣu dataset. Results for all classes of PadChest are consistent and can be found in Appendix 0.A.
On the PadChest dataset, we found that every model is able to obtain good performances when trained and evaluated on similar masking even on images without lungs using the bounding box which removes a high proportion of the image (AUC between 0.85 and 0.93 on the effusion class and minimum 0.74 on the pneumonia with models trained on No lungs BB). These values are all above the baseline results obtained with the logistic regression classifiers on tabular data (mean AUC: 0.71 atelectasis, 0.77 cardiomegaly, 0.72 effusion, 0.60 pneumonia, 0.70 pneumothorax), it is however interesting to note that these AUCs are also above the random value of 0.5. Another interesting result is that for all the classes, models trained on Full images performed better when evaluated on images without lungs than on images with only lungs. These results show that the models are less able to distinguish images when keeping only the lungs with the precise masks with even some classes near or below the random value (atelectasis, effusion, and pneumonia). However, we see an important increase when using the bounding box which may indicate the usage of elements near the lungs captured when using the bounding box. It is worth noting that depending on the receptive field of the model, features at the boundaries of the lungs in Only lungs images may be impacted by the masking.
Fig 5 shows the evolution of the AUC when increasing the mask size for images with only the lungs and without lungs applied to the model trained of Full images. Fig. 18 in Appendix 0.B shows examples of images obtained at each dilation factor. For most classes, the highest drop of performance for images without lungs happens at a dilation factor of 100 or higher when most of the image is masked out. It is interesting to note that despite a drop, the AUC for the pneumonia class remains above random until completely masking out the images. For images with only the lungs we can see that a dilation factor of 50 is needed before achieving results close to Full images. As we concluded from the AUCs using the bounding boxes, these results may reveal the usage of elements near the lungs but also sometimes at the border of the image.
The combination of these results, consistent over all classes, reinforced our hypothesis of models affected by spurious correlation in the dataset.
(a) Mean AUC for the effusion class in PadChest
(b) Mean AUC for the glaucoma class in Chákṣu
Figure 4: Mean AUC across the five models from 5-fold cross-validation on the testing set with different masking for training and evaluation images. The x-axis shows the masking strategy of the images used to evaluate the model, the y-axis shows the masking strategy of the images used to train the model. The color range in both figures is different to fit the range of the specific set.
(a) Images without lungs
(b) Images with only lungs
Figure 5: Evolution of the AUC with standard deviation of the models trained on Full images applied to (a) images with only lungs and (b) images without the lungs while dilating the mask.
On the Chákṣu dataset, we can see random or near-random results on images without the region of interest with an 0.58 AUC for models trained and evaluated on No disc and 0.44 for models trained and evaluated on No disc BB. On the other hand, we obtain better performances for Full images (0.68 AUC), Only disc BB (0.64 AUC) and Only disc (0.78 AUC). These performances are more aligned with what we would expect from models not affected by shortcut learning.
To further test whether models could suffer from shortcut learning, we also assess how models are affected by the optic disc size. In practice, the cup-to-disc ratio or the rim-to-disc ratio are indicators of glaucoma but the size of the disc alone is not [58].
We evaluate the evolution of AUC when increasing the size of the masks for both Only disc and No disc. In this experiment, we either only increase the size of the masks for healthy images or glaucoma images. Fig. 19 in Appendix 0.B shows example of both type of image at different dilation factor.
The results in Fig. 6 show that disk size impacts the models’ outputs. For the No disc model, the AUC increases for larger masks of glaucoma images until a dilation factor of 150 when it starts decreasing. For the model trained on Only disc, the AUC reaches 1.0 at the smallest dilation factor of 5 and never decreases even at the highest value of 500. Note that this seems to also be a problem among clinicians as smaller discs may be more misclassified as healthy and larger discs as glaucoma [59]. Moreover, as the optic disc size varies across populations [60], such spurious correlation could therefore lead to degraded performance in underrepresented populations.
(a) Trained on images without the optic disc (No disc)
(b) Trained on images without the outside of the optic disc (Only disc)
Figure 6: Evolution of the AUC with standard deviation when increasing mask size of healthy images (in blue) and glaucoma images (in orange)
4.2 Poor out-of-distribution performance
We hypothesised that the models trained without the region of interest would have lower performances on external datasets as shortcuts are often dataset-specific, relying on such shortcuts would therefore not transfer well to other datasets. Results in Table 2 seem to partly align with our intuition. We indeed find poor performances of models trained without the ROI, almost always close to an AUC of 0.5. However, we can see that the results of models trained with only the ROI are often also close to this value. For cardiomegaly, effusion, pneumonia and glaucoma, the best performances are achieved by models trained on Full images. Moreover, we observe no statistically significant difference between Full and Only lung on atelectasis for 3/5 folds. However, the results remain well below the results on the training dataset. This result confirms the poor transferability across populations and the need to test and fine-tune models on the target population before applying the models in a different context like a different hospital.
Masking Atelectasis Cardiomegaly Effusion Pneumonia Pneumothorax Glaucoma
No ROI BB 0.53+/-0.004 0.54+/-0.026 0.58+/-0.005 0.52+/-0.009 0.54+/-0.016 0.49+/-0.045
No ROI 0.52+/-0.005 0.54+/-0.022 0.63+/-0.004 0.55+/-0.018 0.51+/-0.020 0.52+/-0.050
Full 0.59+/-0.004 0.72+/-0.016* 0.72+/-0.002* 0.62+/-0.012* 0.56+/-0.016 0.65+/-0.048*
Only ROI BB 0.60+/-0.002 0.67+/-0.012 0.67+/-0.009 0.59+/-0.011 0.51+/-0.011 0.52+/-0.030
Only ROI 0.54+/-0.013 0.58+/-0.022 0.53+/-0.021 0.55+/-0.014 0.60+/-0.029* 0.53+/-0.031
Table 2: Out-of-distribution evaluation results - Mean AUC (+/- standard deviation) of 5-fold models per masking type and class on full images of the ChestX-ray14 dataset’s test set (Atelectasis, Cardiomegaly, Effusion, Pneumonia and Pneumothorax) and AIROGS dataset (Glaucoma). * indicates statistical significance for three or more folds (p-values < 0.05)
4.3 Closer embeddings between full images and without ROI
We show the results from the t-SNE algorithm on the PadChest dataset in Fig. 7(a). We find that the embeddings of each masking type clearly cluster together with a small overlapping between embeddings of the full images, images without lungs and images with only lungs using the bounding box. On the Chákṣu dataset in Fig. 7(b), we see the embeddings of Full images, No disc and No disc BB completely overlapping while Only disc and Only disc BB have a clear cluster.
For chest X-rays, the mean cosine similarity in Table 3 is higher when comparing full images with images without lungs using masks, followed by images with only the lungs using the bounding box, then images without lungs using the bounding box and finally, the lowest similarity is with images without lungs using masks. Both similarities of No lungs and Only lungs BB are close, which could also confirm the usage of elements at the border or between the lungs that are not included in the images of Only lungs using the mask.
For eye fundus, the results in Table 4 are aligned with the t-SNE and show the high proximity of images without the ROI and Full images compared to images with only the ROI. It however doesn’t align with the AUC measure in Fig. 4(b) as we would expect AUC of Full images to be closer to No ROI images than Only ROI images. The close embeddings before the classification head may be caused by the similarity of base images as Full and No lungs images share more common pixels than Full and Only lungs.
These results show that despite being commonly used, the interpretation of t-SNE visualisations and cosine similarities of features before the classification head is limited, should be considered carefully, and completed with further analysis.
(a) PadChest
(b) Chákṣu
Figure 7: t-SNE of the embeddings before classification head of the model trained with the full images dataset on images with various masking from (a) PadChest dataset and (b) the Chákṣu dataset. In blue circles are full images, in orange triangles without ROI using the mask, in green squares without the ROI using the bounding box, in purple stars with only the ROI using the bounding box and in red crosses only the ROI using the mask
Image label All Cardiomegaly Pneumonia Atelectasis Pneumothorax Effusion
No lungs 0.90+/-0.02 0.90+/-0.02 0.90+/-0.02 0.90+/-0.02 0.91+/-0.02 0.90+/-0.02
No lungs BB 0.83+/-0.04 0.83+/-0.04 0.85+/-0.04 0.86+/-0.04 0.85+/-0.04 0.85+/-0.04
Only lungs 0.76+/-0.03 0.74+/-0.03 0.76+/-0.04 0.75+/-0.03 0.79+/-0.02 0.73+/-0.03
Only lungs BB 0.89+/-0.03 0.88+/-0.03 0.88+/-0.04 0.87+/-0.04 0.89+/-0.04 0.86+/-0.04
Table 3: Mean cosine similarity (+/- standard deviation) over pairs of images of the first fold, between embeddings of images with different masking by the model trained on full images of the PadChest dataset
Image label All Glaucoma
No disc 0.97+/- 0.01 0.97+/- 0.01
No disc BB 0.98+/- 0.01 0.98+/- 0.01
Only disc 0.69+/- 0.03 0.71+/- 0.04
Only disc BB 0.68+/- 0.03 0.69+/- 0.03
Table 4: Mean cosine similarity (+/- standard deviation) over images of the first fold, between embeddings of images with different masking by the model trained on full images of the Chákṣu dataset
4.4 Models can use both the relevant structure and the shortcut
Fig. 8 shows examples of behaviours found with the explainability maps from SHAP on the PadChest dataset. We can for example see the usage of pacemakers in the decision-making for the cardiomegaly class (Fig. 8(a)) even though the model still uses the relevant part of the image (i.e. the heart) in the decision-making (Fig. 8(b)). It is interesting to note that while there can be a correlation between cardiomegaly and pacemakers there is no causal link between the two. Similarly to tubes used as shortcuts for pneumothorax classification [5, 6], this shows an example of shortcuts that can be easily detected and are medically understandable. On the other hand, explainability maps of images without the ROI highlight the usage of elements at the edge of the image even far from the region of interest (e.g. the heart for cardiomegaly such as in Fig. 8(c)). In this case, the regions highlighted by SHAP don’t present any apparent reason for their usage. It is therefore difficult to provide a clear explanation of the model output. Moreover, SHAP provides a local explanation of the output and may therefore not be able to detect possible global shortcuts such as specific noise created by different scanners or settings during the acquisition. These results highlight an important distinction between the non-obvious shortcuts, such as noise, and the medically understandable shortcuts, such as pacemakers, that can be more easily detected and therefore mitigated.
(a) With pacemaker
(b) Without pacemaker
(c) Without the lungs
Figure 8: SHAP values of the cardiomegaly class with and without a pacemaker, and without the lungs region. Red values indicate parts increasing the probability. The blue arrows in (a) show the pacemaker location.
For the Chákṣu dataset, we can also see in Fig. 9 the usage of eye boundaries in the decision, these areas contain differences for the 3 cameras with for example the text in the Remedio images or the cropped boundaries in the Forus images. SHAP also highlights the optic disc but the finding in Section 4.1 put doubt on the correct usage of this area by the model as it could also be only looking at the disc size and not the cup-to-disc ratio. These two results highlight the importance of not only relying on explainability methods to evaluate models’ behaviour but also conducting more thorough analyses.
Figure 9: SHAP values of an image with glaucoma
4.5 Lack of ROI and context makes the diagnosis difficult for a domain expert
Out of the 35 cases of the conditions present in images without the ROI (using the mask or bounding box) only 2 were found by the radiology resident, both in images with lungs removed using the mask. One was a patient with cardiomegaly which is diagnosed using the heart still present when only removing the lungs. This result strengthens our hypothesis that the model relies on non-relevant features to domain experts. Results on full images and images containing only the ROI also reveal the difficulty of the task when only relying on the image. A recurring observation across images is the lack of information about the projection as in Fig. 10.
Due to the automatic selection of images based on probabilities, we also found cases in Fig. 11 of quality problems both in the original and the mask dataset. This highlights one of the problems of using large-scale public datasets for which the assessment of quality is difficult but essential as shown on other tasks and datasets to obtain better and more robust models [61, 62].
Figure 10: One of the images labelled by the radiologist with the comment "The diagnosis would be heavily dependent on whether it is AP or PA. Also potential effusion would be hard to exclude on AP"
(a) Image with poor quality.
(b) Image with a non-accurate mask.
Figure 11: Example of images with either a bad quality making the diagnosis difficult for a radiologist due to the blur in the whole image or with a non-accurate mask indicated by the red box outside the chest area.
5 Discussion
We studied how convolutional neural networks are affected by the masking of different parts of medical images to highlight the effect of shortcut learning. We analyzed performances, comparing the AUC of each masking strategy, but also went beyond by using an explainability method and comparing the embeddings. Our experiments showed the ability of CNN models to rely on non-relevant parts of medical images. Our results highlight the difficulty of detecting such behaviours. The AUC obtained on different masking configurations may show strong evidence of shortcut learning, like in chest X-rays. However, it can also be more hidden like in glaucoma detection for which the performance analysis reveals no signs of shortcut learning. With eye-fundus photography, we also showcased a limitation of relying on the explainability method as the model could be focusing on a correct region but not using the correct feature. It also shows the potential difficulty of detecting patient-level shortcuts described in [34] such as anatomical confounders when they are located at the area of interest.
The results from the study with the radiology resident highlight how relying on a single source of information (e.g. an image) makes the diagnosis more difficult or even impossible in some cases. In this setting, the lack of context may guide the model to rely more on shortcuts (such as pacemakers). The recent development of multimodal models [63] may therefore help mitigate these issues as the model has access to more clinically relevant information, although it is yet unknown whether the additional modalities could introduce other shortcuts. Importantly, for certain pathologies, chest X-rays are only one part of the diagnostic process, for example a chest X-ray could be used for triage before a follow-up CT scan.
We only used a single CNN architecture, a Densenet-121, for our experiments. Other works showed the effect of shortcut learning on other CNN architectures in medical imaging [26, 64]. Other architectures such as transformers or vision-language models are becoming popular in medical image analysis, but their robustness to shortcuts has only been studied in the natural images domain [65]. As our code repository allows us to easily adapt the framework to other models, it would be of interest to apply it to these different types of architectures in future works.
We focused on the discriminative capacity of models using the AUC as the metric. However, other aspects, such as calibration or fairness across different subgroups, need to be taken into account during the evaluation, especially if the aim is to deploy the model in a clinical environment [66]. Finally, our method only provides a way to evaluate the risk of shortcut learning, further analysis may be performed to conclude on its absence or presence and other techniques for improving data quality and/or mitigating the model bias would therefore need to be applied.
We believe that the analysis of potential biases should be an important step in the evaluation of AI models. In addition to the poor transferability to other domains such as another hospital, these spurious correlations can have a major impact on the diagnosis of underrepresented populations. We also believe that our field would benefit from going beyond widely used performance metrics such as the AUC that don’t necessarily translate into real clinical progress. To this aim, collaboration with hospitals and domain experts is essential to ground research in more clinical knowledge, ensuring better data understanding but also the relevancy and added value of developed algorithms from a clinical point of view.
Both our and other recent works show the need to question whether models the community trains actually learn salient characteristics of the diseases. It is likely that published models overfit to spurious features, and should not be applied in practice until the community figures out how to resolve these issues and achieve improved robustness.
5.0.1 Acknowledgments
We thank Casper Anton Poulsen, Trine Naja Eriksen and Cathrine Damgaard for their early work within this research line. We also thank Evangelia Christodoulou for her advice on the statistical significance test. This work was supported by the DFF - Inge Lehmann 1134-00017B and Novo Nordisk Foundation NNF21OC0068816.
Appendix 0.A AUC per label and masking
(a) on 5-folds validation sets
(b) on test set
Figure 12: Atelectasis mean AUC across 5-fold models with different masking for training and evaluation images
(a) on 5-folds validation sets
(b) on test set
Figure 13: Cardiomegaly mean AUC across 5-fold models with different masking for training and evaluation images
(a) on 5-folds validation sets
(b) on test set
Figure 14: Effusion mean AUC across 5-fold models with different masking for training and evaluation images
(a) on 5-folds validation sets
(b) on test set
Figure 15: Pneumonia mean AUC across 5-fold models with different masking for training and evaluation images
(a) on 5-folds validation sets
(b) on test set
Figure 16: Pneumothorax mean AUC across 5-fold models with different masking for training and evaluation images
(a) on 5-folds validation sets
(b) on test set
Figure 17: Glaucoma mean AUC across 5-fold models with different masking for training and evaluation images
Appendix 0.B Example of images with different dilation factors
(a) Images without lungs
(b) Images with only lungs
Figure 18: Example of NoLungs and OnlyLungs images while increasing the mask size. In this example, the mask covers the entire image at a dilation factor of 300 or more, but a higher dilation may be needed depending on the original mask.
(a) Images without the optic disc
(b) Images with only the optic disc
Figure 19: Example of NoDisc and OnlyDisc images while increasing the mask size.